OpenStudy (anonymous):

Given: 1/3x + y = 15. Which of the following are ordered pairs of this equation? {(-3, 14),(0, 15),(3, 16)} {(-3, 16),(0, 15),(3, 14)} {(16, -3),(15, 0),(14, 3)}

OpenStudy (ccswims):

first get it to y=mx + b so subtract -1/3x to both sides y = -1/3x + 15 now just graph it a see what coordinates they gave are on the line

OpenStudy (anonymous):

Ok, thanks, @ccswims :)